# AERO-MPPI: Anchor-Guided Ensemble Trajectory Optimization for Agile Mapless Drone Navigation

### v1

@@ Abstract
Agile mapless navigation in cluttered 3D environments poses significant challenges for autonomous drones. Conventional mapping–planning–control pipelines incur high computational cost and propagate estimation errors. We present AERO-MPPI, a fully GPU-accelerated framework that unifies perception and planning through an anchor-guided ensemble of Model Predictive Path Integral (MPPI) optimizers. Specifically, we design a multi-resolution LiDAR point-cloud representation that rapidly extracts spatially distributed “anchors” as look-ahead intermediate endpoints, from which we construct polynomial trajectory guides to explore distinct homotopy path classes. At each planning step, we run multiple MPPI instances in parallel and evaluate them with a two-stage multi-objective cost that balances collision avoidance and goal reaching. Implemented entirely with NVIDIA Warp GPU kernels, AERO-MPPI achieves real-time onboard operation and mitigates the local-minima failures of single-MPPI approaches. Extensive simulations in forests, verticals, and inclines demonstrate sustained reliable flight above 7 m/s, with success rates above 80% and smoother trajectories compared to state-of-the-art baselines. Real-world experiments on a LiDAR-equipped quadrotor with NVIDIA Jetson Orin NX 16G confirm that AERO-MPPI runs in real time onboard and consistently achieves safe, agile, and robust flight in complex cluttered environments. The code will be open-sourced upon acceptance of the paper.
@@ I Introduction
Mapless navigation generates and executes collision-free motions using only local, onboard perception (e.g., LiDAR point clouds). This avoids the error accumulation inherent in maintaining a global map and thereby enhances flight agility. Among sampling-based trajectory optimizers, Model Predictive Path Integral (MPPI) control [5] has emerged as a strong candidate because it flexibly accommodates diverse objective functions, system dynamics, and constraints. More importantly, its sampling-based optimization is well suited to parallelization.
@@
  • We conducted extensive validation in simulation across three challenging scenarios. Results show that our method achieves improved agility, success rate, and smoothness compared with strong baseline planners [3, 2, 1, 4], while maintaining real-time performance, safety, and deployability.
@@ II RELATED WORK
The aforementioned works rely on gradient information from explicit maps. In contrast, MPPI control [5] is a sampling-based, model-predictive method that repeatedly draws random control sequences, simulates (rolls out) the system’s nonlinear dynamics under those controls, and evaluates a cost for each trajectory. The controls are then updated—via a path-integral weighted average of the samples, favoring lower cost trajectories and yielding an optimized control action at each planning step. This formulation naturally handles nonlinear dynamics and non-smooth obstacle costs, while GPU parallelization enables thousands of rollouts in real time.
@@ IV-B2 Refining the initial endpoints
$\bm{p}^{\text{ref}}_{I,J}\;=\;\bm{p}_{0}+\ell\cdot\bm{v}_{I,J},$ (8)
where $\bm{v}_{I,J}$ is computed as in (7). This enforces that anchors strictly follow the locally safest heading provided by each cell in the coarse partition.
@@ V-C1 Hardware details
We deployed our method on a 380 mm quadrotor platform weighing 1.65 kg, equipped with a Livox Mid-360 LiDAR, a PX4-based embedded autopilot, and a Jetson Orin NX 16 GB onboard computer (see Fig. 7). The PX4 Extended Kalman Filter (EKF) runs at 100 Hz, fusing IMU measurements with Fast-LIO2 localization [17] to provide high-frequency pose estimates. Using these estimates, AERO-MPPI runs on the Jetson’s GPU for real-time inference and replans at 50 Hz to generate control commands that the PX4 autopilot executes.

### v2

@@ Abstract
Agile mapless navigation in cluttered 3D environments poses significant challenges for autonomous drones. Conventional mapping–planning–control pipelines incur high computational cost and propagate estimation errors. We present AERO-MPPI, a fully GPU-accelerated framework that unifies perception and planning through an anchor-guided ensemble of Model Predictive Path Integral (MPPI) optimizers. Specifically, we design a multi-resolution LiDAR point-cloud representation that rapidly extracts spatially distributed “anchors” as look-ahead intermediate endpoints, from which we construct polynomial trajectory guides to explore distinct homotopy path classes. At each planning step, we run multiple MPPI instances in parallel and evaluate them with a two-stage multi-objective cost that balances collision avoidance and goal reaching. Implemented entirely with NVIDIA Warp GPU kernels, AERO-MPPI achieves real-time onboard operation and mitigates the local-minima failures of single-MPPI approaches. Extensive simulations in forests, verticals, and inclines demonstrate sustained reliable flight above 7 m/s, with success rates above 80% and smoother trajectories compared to state-of-the-art baselines. Real-world experiments on a LiDAR-equipped quadrotor with NVIDIA Jetson Orin NX 16G confirm that AERO-MPPI runs in real time onboard and consistently achieves safe, agile, and robust flight in complex cluttered environments. Code is available at https://github.com/XinChen-stars/AERO_MPPI.
@@ I Introduction
Mapless navigation generates and executes collision-free motions using only local, onboard perception (e.g., LiDAR point clouds). This avoids the error accumulation inherent in maintaining a global map and thereby enhances flight agility. Among sampling-based trajectory optimizers, Model Predictive Path Integral (MPPI) control [15] is a strong candidate because it flexibly accommodates diverse objectives, system dynamics, and constraints. More importantly, its sampling-based optimization is well-suited to parallelization.
@@
  • We conducted extensive validation in simulation across challenging scenarios. Results show our method achieves improved agility, success rate, and smoothness compared with baseline planners [21, 19, 20, 7], while maintaining real-time performance, safety, and deployability.
@@ II RELATED WORK
The aforementioned works rely on gradient information from explicit maps. In contrast, MPPI control [15] is a sampling-based, model-predictive method that repeatedly draws random control sequences, simulates (rolls out) the system’s nonlinear dynamics under those controls, and evaluates a cost for each trajectory. The controls are then updated via a path-integral weighted average of the samples, favoring lower cost trajectories and yielding an optimized control action at each planning step. This formulation naturally handles nonlinear dynamics and non-smooth obstacle costs, while GPU parallelization enables thousands of rollouts in real time.
@@ IV-B2 Refining the initial endpoints
$\bm{p}^{\text{ref}}_{I,J}=\bm{p}_{0}+\ell\cdot\bm{d}_{I,J},$ (8)
where $\bm{d}_{I,J}$ is the unit direction vector as defined in (7).
This enforces that anchors strictly follow the locally safest heading provided by each cell in the coarse partition.
@@ V-C1 Hardware details
We deployed our method on a 380 mm quadrotor platform weighing 1.65 kg, equipped with a Livox Mid-360 LiDAR, a PX4-based embedded autopilot, and a Jetson Orin NX 16 GB onboard computer (see Fig. 7). The PX4 Extended Kalman Filter (EKF) runs at 100 Hz, fusing IMU measurements with Fast-LIO2 localization [16] to provide high-frequency pose estimates. Using these estimates, AERO-MPPI replans control sequences at 10 Hz on the Jetson GPU, while the PX4 autopilot executes the commands at 50 Hz in a receding-horizon manner.